OpenStudy (anonymous):

What is the exact value for the expression square root of 171 - square root of 19 + square root of 76 ? Simplify if possible. plzz help

OpenStudy (coconutjj):

171, 19 and 76 all have a GCF of 19, 171/19 = 9 19/19 = 1 76/19 = 4 Now we can conclude: 9 x 19 = 171 1 x 19 = 19 4 x 19 = 76 Now we know\[\sqrt{ab} = \sqrt{a} \times \sqrt{b}\] So \[\sqrt{(9)\times19}=171\] \[\sqrt{(1)\times19}=19\] \[\sqrt{(4)\times19}=76\] Now seperate the square root, so it is sqrt(a) x sqrt(b) Take the square root of the perfect square...(in brackets) so we get \[3\sqrt{19}\] \[1\sqrt{19}\] \[2\sqrt{19}\] Add/Subtract with the appropriate sign... We get \[4\sqrt{19}\] as a result